# Sirui Xu1  Hung Yu Ling2  Yu-Xiong Wang1†  Liang-Yan Gui1†

### v1

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: InterMimic enables physically simulated humans to perform interactions with dynamic and diverse objects. It supports highly-dynamic, multi-object interactions and scalable skill learning (Top), making it adaptable for versatile downstream applications (Bottom): it can translate whole-body loco-manipulation skills to a humanoid robot [130, 47], perfect interaction MoCap data, and bridge kinematic generation, e.g., predicting future interactions from past (InterDiff [160]) or generating interactions given text prompts (InterDreamer [162]).
[…]
.tocmtchapter
@@ 1 Introduction
Introducing teacher policies offers several key benefits. By leveraging teacher rollouts, we effectively distill raw MoCap data into refined HOI references with a unified embodiment and enhanced physical fidelity. These refined references guide the subsequent student policy training, reducing the negative impact of errors in the original MoCap data. A major hurdle in scaling motion imitation is the sample inefficiency of Reinforcement Learning (RL), which can lead to prohibitively long training times. Our teacher-student approach mitigates this through a space-time trade-off: multiple teacher policies are trained in parallel on smaller, more manageable data subsets, and their expertise is then distilled into a single student policy. We begin with demonstration-based distillation to bootstrap PPO [114] updates, reducing reliance on pure trial and error and enabling more effective scaling. As training progresses, the student gradually shifts from heavy demonstration guidance to increased RL updates, ultimately surpassing simple demonstration memorization. This mirrors alignment strategies in Large Language Models (LLMs), where demonstration-based pretraining is refined through RL fine-tuning [98, 129].
To summarize, our contributions are as follows: (i) We introduce InterMimic, which, to the best of our knowledge, is the first framework designed to train physically simulated humans to develop a wide range of whole-body motor skills for interacting with diverse and dynamic objects, extending beyond traditional grasping tasks. (ii) We develop a teacher-student training strategy, where teacher policies provide a unified solution to address the challenges of retargeting and refining in HOI imitation. The student distillation introduces a scalable solution by leveraging a space-time trade-off. (iii) We demonstrate that our unified framework, InterMimic, as illustrated in Figure 1, effectively handles versatile physics-based interaction animation, recovering motions with realistic and physically plausible details. Notably, by combining kinematic generators with InterMimic, we enable a physics-based agent to achieve tasks such as interaction prediction and text-to-interaction generation.
@@ 2 Related Work
Significant progress has been made in human interaction animation and control, with advancements in areas such as human-human interactions [161, 31, 81, 144, 158, 84, 44, 55, 75], hand-object interactions [121, 145, 169, 178, 170, 195, 197, 7, 183, 94, 128, 14, 101, 128, 1, 168, 139, 159], single-frame interactions [153, 184, 141, 107, 43, 60, 163, 164, 20, 73, 166, 56, 167, 165, 181, 190, 192, 149], human interactions with static scenes [36, 136, 137, 9, 99, 152, 125, 65, 140, 69, 188, 85, 57, 16, 185, 8, 95, 193], and real-world humanoid control for object manipulation [33, 115, 38, 27, 28, 11, 175, 12, 135, 52, 86, 177, 39, 15, 4, 25, 72, 54, 48, 23, 76]. In this section, we focus on recent studies on whole-body interaction animation, particularly involving dynamic objects.
@@ 2.1 Kinematic Interaction Animation
Generating human interactions has been a long-standing topic in animation and computer graphics [66, 32]. Significant advances in character animation [10, 196, 21, 155, 68, 77, 194, 3] have emerged with the advent of deep learning, e.g., including phase-function-based methods [41] that enable object interactions like carrying a box [119] or playing basketball [120]. This is extended to more diverse but static objects approaching [186, 150, 123, 64]. Subsequent efforts [30, 109, 71, 151, 74, 51, 50, 87] integrate object motion into interactions but remain constrained by assuming that interactions occur primarily through the hands. To address this, recent developments [17, 134, 160, 103, 24, 148, 22, 118, 58, 162, 40] introduce interactions in a fashion of whole-body loco-manipulation that engages multiple body parts in contact. However, these methods often suffer from physical inaccuracies, such as floating contacts and penetrations, while they generate only body motion without considering hand motion [88] or dexterity. In this work, we address physical inaccuracies by refining imperfect kinematic generation through physics simulation, with InterDiff [160] and HOI-Diff [103] serving as motion planning for loco-manipulation that bridges high-level decision-making (e.g., text instruction) with low-level execution.
@@ 2.2 Physics-based Interaction Animation
Physics-based methods generate motion through motor control policies within a physics simulator, e.g., achieved via deep reinforcement learning to track reference motions [104]. These policies are directly applicable for executing simple interactions, such as punching or striking an object [106, 124, 18, 127]. To achieve more complex interactions, early studies focus on specific scenarios, including notable sports-related [92] examples such as basketball [142, 79, 143], skating [78], soccer [156, 42, 80, 35, 18], tennis [179], table tennis [138], as well as tasks proposed in [2]. Research also demonstrates flexibility in more general but simpler box carrying tasks [189]. These advancements are achieved through the integration of multiple control policies [97], the use of adversarial motion priors [37, 105, 29], and imitating diverse kinematic generations [157, 151]. However, these methods train their policies in a non-scalable manner, with each policy handling only specific object types or actions. In pursuit of a single, scalable policy to enable multiple interaction skills, existing methods either rely on fixed interaction patterns, such as approaching and grasping objects by following predefined trajectories [6], or remain confined to single-hand grasping actions [91]. Additionally, they depend on highly curated data from the GRAB dataset [122], which, despite its high quality, primarily features low-dynamic full-body motion and only small-sized objects. More recently proposed datasets [5, 49, 46, 180, 26, 70, 191, 59, 93, 176, 154, 182, 187, 82, 83] offer richer full-body interactions with objects in diverse shapes but contain significant artifacts that challenge existing motion imitation approaches. We process data from OMOMO [70], BEHAVE [5], HODome [180], IMHD [191], and HIMO [93] into the simulator, demonstrating InterMimic’s scalability for diverse interactions and robustness to MoCap artifacts.
@@ 3 Methodology
Task Formulation. The goal of human-object interaction (HOI) imitation is to learn a policy $\pi$ that produces simulated human-object motion $\{\boldsymbol{q}_{t}\}_{t=1}^{T}$ closely matching a ground-truth reference $\{\hat{\boldsymbol{q}}_{t}\}_{t=1}^{T}$ derived from large-scale MoCap data. Given the geometries of the human and objects, the policy should also compensate for missing or inaccurate details in the dataset. Each pose $\boldsymbol{q}_{t}$ has two components: the human pose $\boldsymbol{q}^{h}_{t}$ and the object pose $\boldsymbol{q}^{o}_{t}$. The human pose is defined as $\boldsymbol{q}^{h}_{t}=\{\boldsymbol{\theta}^{h}_{t},\boldsymbol{p}^{h}_{t}\}$, where $\boldsymbol{\theta}^{h}_{t}\in\mathbb{R}^{51\times 3}$ represents the joint rotations, and $\boldsymbol{p}^{h}_{t}\in\mathbb{R}^{51\times 3}$ specifies the joint positions. Specifically, our human model includes 30 hand joints and 21 joints for the rest of the body. The object pose $\boldsymbol{q}^{o}_{t}$ is represented as $\{\boldsymbol{\theta}^{o}_{t},\boldsymbol{p}^{o}_{t}\}$, where $\boldsymbol{\theta}^{o}_{t}\in\mathbb{R}^{3}$ denotes the object’s orientation and $\boldsymbol{p}^{o}_{t}\in\mathbb{R}^{3}$ the position. All simulation states have corresponding ground-truth values, denoted by the hat symbol. For instance, the reference object rotation is $\{\hat{\boldsymbol{\theta}}^{o}_{t}\}_{t=1}^{T}$. The environmental setup for the simulation is detailed in Sec. B of supplementary.
@@ 3.1 Policy Representation
State. The state $\boldsymbol{s}_{t}$, which serves as input to the policy, comprises two components $\boldsymbol{s}_{t}=\{\boldsymbol{s}_{t}^{s},\boldsymbol{s}_{t}^{g}\}$. The first part, $\boldsymbol{s}_{t}^{s}$, contains human proprioception and object observations, expressed as, $\{\{\boldsymbol{\theta}_{t}^{h},\boldsymbol{p}_{t}^{h},\boldsymbol{\omega}_{t}^{h},\boldsymbol{v}_{t}^{h}\},\{\boldsymbol{\theta}_{t}^{o},\boldsymbol{p}_{t}^{o},\boldsymbol{\omega}_{t}^{o},\boldsymbol{v}_{t}^{o}\},\{\boldsymbol{d}_{t},\boldsymbol{c}_{t}\}\},$ where $\{\boldsymbol{\theta}_{t}^{h},\boldsymbol{p}_{t}^{h},\boldsymbol{\omega}_{t}^{h},\boldsymbol{v}_{t}^{h}\}$ represent the rotation, position, angular velocity, and velocity of all joints, respectively, while $\{{\boldsymbol{\theta}}_{t}^{o},\boldsymbol{p}_{t}^{o},{\boldsymbol{\omega}}_{t}^{o},\boldsymbol{v}_{t}^{o}\}$ represent the orientation, location, velocity, and angular velocity of the object, respectively. Motivated by [13], we include object geometry and whole-body haptic sensing from two elements: (i) $\boldsymbol{d}_{t}$, vectors from human joints to their nearest points on each object surface; and (ii) $\boldsymbol{c}_{t}$, contact markers indicating whether the human’s rigid body parts experience applied forces; this serves as simplified tactile or force sensing – an important multi-modal input in robot manipulation tasks [19, 171, 116, 45]. The goal state $\boldsymbol{s}_{t}^{g}=\{\boldsymbol{s}_{t,t+k}^{g}\}_{k\in K}$ integrates reference poses from the ground truth motion, where $\boldsymbol{s}_{t,t+k}^{g}$ is defined as,
@@
We extract reference contact markers $\hat{\boldsymbol{c}}_{t+k}$ by inferring dynamic information, in addition to inaccurate contact distances, specifically by analyzing the object’s acceleration to detect human-induced forces. To accommodate the variability in contact distances observed in reference motion, we discretize reference contact markers using varying distance thresholds, as illustrated in Fig. 3(i). The neutral areas serve as buffer zones, avoiding the penalization or enforcement of strict contact. See Sec. C of supplementary for details.
@@ 3.2 Imitation as Perfecting
Imitation as Retargeting. We tailor teacher policies to each human subject, while all policies share the same base human model. This serves the retargeting purpose by converting HOIs from different human shapes into a unified base shape. Although motion imitation does not necessarily require a unified human model [146, 89], our approach offers two benefits: (i) It enhances integration with kinematic generation methods, which generally perform better on a single, unified shape [34, 126]. (ii) It demonstrates possible integration with real-world humanoid deployment, which requires retargeting to a consistent physical embodiment. In Figure 1, our method translates MoCap data into motor skills on a Unitree G1 [130] with two Inspire hands [47], all without external retargeting in complex contact-rich scenarios. See Sec. F of the supplementary for additional details.
@@
[…]
@@
Physical State Initialization. Learning later-phase motion can be essential for policies to achieve high rewards during earlier phases, compared to incrementally learning from the starting phase. Thus, Reference State Initialization (RSI) [104] sets the current pose ${\boldsymbol{q}}_{t}$ to a reference pose $\hat{\boldsymbol{q}}_{t}$ at a random timestep $t$, for initializing the rollout. However, initializing with the imperfect reference can introduce critical artifacts, such as contact floating or incorrect hand motion, leading to unrecoverable failures, e.g., object falling, as depicted in Figure 3(ii). These issues render many initializations ineffective, limiting training on certain interaction phases since successful rollouts may not reach them before the maximum length. The problem is exacerbated by the use of prioritized sampling [90, 156, 146, 125], which favors high-failure-rate initializations.
@@ 3.3 Imitation with Distillation
Reference Distillation. Noisy MoCap data can hinder policy learning, especially at larger scales. In contrast, teacher policies trained on smaller-scale data effectively address these issues by correcting contact artifacts, refining hand placements, and recovering missing details (see Figures 1 and 5). To fully leverage teacher policies, we use their rollouts as references for defining the student policy’s goal state and reward functions, distinguishing our approach from distillation based on only action output.
Policy Distillation. We also apply distillation on action outputs, which we view as crucial for scaling policies to large datasets. In essence, we trade space for time: teacher policies are trained in parallel on smaller data subsets, allowing the student policy to scale through distillation. Following Algorithm 1, we begin with Behavior Cloning (BC) [53, 133] and then use RL fine-tuning to go beyond demonstration memorization, an approach common in LLM alignment [98, 129]. Inspired by [108, 117], we integrate BC into online policy updates and adopt a staged schedule: we start with DAgger [112] and gradually transition to PPO. Throughout, the critic is continuously trained with the reward from Sec. 3.2. This RL fine-tuning phase is crucial because teacher policies may behave differently when performing similar skills, and simple BC can lead to suboptimal “averaging” behavior, where RL fine-tuning helps the student policy converge on optimal solutions.

### v2

@@ 1 Introduction
Introducing teacher policies offers several key benefits. By leveraging teacher rollouts, we effectively distill raw MoCap data into refined HOI references with a unified embodiment and enhanced physical fidelity. These refined references guide the subsequent student policy training, reducing the negative impact of errors in the original MoCap data. A major hurdle in scaling motion imitation is the sample inefficiency of Reinforcement Learning (RL), which can lead to prohibitively long training times. Our teacher-student approach mitigates this through a space-time trade-off: multiple teacher policies are trained in parallel on smaller, more manageable data subsets, and their expertise is then distilled into a single student policy. We begin with demonstration-based distillation to bootstrap PPO [71] updates, reducing reliance on pure trial and error and enabling more effective scaling. As training progresses, the student gradually shifts from heavy demonstration guidance to increased RL updates, ultimately surpassing simple demonstration memorization. This mirrors alignment strategies in Large Language Models (LLMs), where demonstration-based pretraining is refined through RL fine-tuning [56].
To summarize, our contributions are as follows: (i) We introduce InterMimic, which, to the best of our knowledge, is the first framework designed to train physically simulated humans to develop a wide range of whole-body motor skills for interacting with diverse and dynamic objects, extending beyond traditional grasping tasks. (ii) We develop a teacher-student training strategy, where teacher policies provide a unified solution to address the challenges of retargeting and refining in HOI imitation. The student distillation introduces a scalable solution by leveraging a space-time trade-off. (iii) We demonstrate that our unified framework, InterMimic, as illustrated in Figure LABEL:fig:teaser, effectively handles versatile physics-based interaction animation, recovering motions with realistic and physically plausible details. Notably, by combining kinematic generators with InterMimic, we enable a physics-based agent to achieve tasks such as interaction prediction and text-to-interaction generation.
@@ 2 Related Work
Significant progress has been made in physics-based human interaction animation and control, with advancements in areas such as human-human interactions [48, 90], hand-object interactions [60, 104, 86, 100], human interactions with static scenes [5, 57, 96, 78, 36], and real-world humanoid control for object manipulation [16, 72, 19, 107, 28, 20, 2, 41, 11, 43]. Among these areas, we are the first to achieve universal whole-body loco-manipulation simulation with diverse dynamic objects, beyond pick-and-place and grasping actions – a novel achievement in animation and an idealized reference for real-world humanoid control. Below, we elaborate on recent studies on whole-body interaction animation, particularly involving dynamic objects.
@@ 2.1 Kinematic Interaction Animation
Generating human interactions has been a long-standing topic in animation and computer graphics [37, 15]. Significant advances in character animation have emerged with the advent of deep learning, e.g., including phase-function-based methods [22] that enable object interactions like carrying a box [74] or playing basketball [75]. This is extended to more diverse but static objects approaching [111, 94, 76, 35]. Subsequent efforts [14, 66, 39, 95, 42, 27, 25, 49] integrate object motion into interactions but remain constrained by assuming that interactions occur primarily through the hands. To address this, recent developments [7, 84, 102, 62, 12, 93, 10, 73, 30, 103, 21] introduce interactions in a fashion of whole-body loco-manipulation that engages multiple body parts in contact. However, these methods often suffer from physical inaccuracies, such as floating contacts and penetrations, while they generate only body motion without considering hand dexterity. In this work, we address physical inaccuracies by refining imperfect kinematic generation through physics simulation, with InterDiff [102] and HOI-Diff [62] serving as motion planning for loco-manipulation that bridges high-level decision-making (e.g., text instruction) with low-level execution.
@@ 2.2 Physics-based Interaction Animation
Physics-based methods generate motion through motor control policies within a physics simulator, e.g., achieved via deep reinforcement learning to track reference motions [63]. These policies are directly applicable for executing simple interactions, such as punching or striking an object [64, 79, 8, 80]. To achieve more complex interactions, early studies focus on specific scenarios, including notable sports-related [52] examples such as basketball [88], skating [44], soccer [98], tennis [108], table tennis [85], and more proposed in [1]. Research also demonstrates flexibility in more general but simpler box carrying tasks [113, 58, 87]. These advancements are achieved through the integration of multiple control policies [55], the use of adversarial motion priors [18, 65, 13], and imitating diverse kinematic generations [99, 95]. However, these methods train their policies in a non-scalable manner, with each policy handling only specific object types or actions. In pursuit of a single, scalable policy to enable multiple interaction skills, existing methods either rely on fixed interaction patterns, such as approaching and grasping objects [4, 50], or extend single-object skills, e.g., interactions involving a basketball [89]. Additionally, they mostly depend on highly curated data from the GRAB dataset [77], which, despite its high quality, primarily features low-dynamic full-body motion and only small-sized objects. More recent datasets [3, 26, 23, 109, 40, 114, 31, 53, 47, 97, 110, 112, 45, 46, 101] offer richer full-body interactions with diverse objects but contain noticeable artifacts that challenge existing motion imitation approaches. We process data from OMOMO [40], BEHAVE [3], HODome [109], and IMHD [114] collected in the InterAct [101] dataset, and the multi-object dataset HIMO [53], highlighting InterMimic’s scalability to diverse interactions and its robustness against MoCap artifacts.
@@ 3 Methodology
Task Formulation. The goal of human-object interaction (HOI) imitation is to learn a policy $\pi$ that produces simulated human-object motion $\{\boldsymbol{q}_{t}\}_{t=1}^{T}$ closely matching a ground-truth reference $\{\hat{\boldsymbol{q}}_{t}\}_{t=1}^{T}$ derived from large-scale MoCap data. Given the geometries of the human and objects, the policy should also compensate for missing or inaccurate details in the dataset. Each pose $\boldsymbol{q}_{t}$ has two components: the human pose $\boldsymbol{q}^{h}_{t}$ and the object pose $\boldsymbol{q}^{o}_{t}$. The human pose is defined as $\boldsymbol{q}^{h}_{t}=\{\boldsymbol{\theta}^{h}_{t},\boldsymbol{p}^{h}_{t}\}$, where $\boldsymbol{\theta}^{h}_{t}\in\mathbb{R}^{52\times 3}$ represents the joint rotations, and $\boldsymbol{p}^{h}_{t}\in\mathbb{R}^{52\times 3}$ specifies the joint positions. Specifically, our human model includes 30 hand joints and 22 joints for the rest of the body, with one root joint’s rotation and position specified in global coordinates, whereas the rotations and positions of all other joints are defined relative to their respective parent joints’ coordinate systems. The object pose $\boldsymbol{q}^{o}_{t}$ is represented as $\{\boldsymbol{\theta}^{o}_{t},\boldsymbol{p}^{o}_{t}\}$, where $\boldsymbol{\theta}^{o}_{t}\in\mathbb{R}^{3}$ denotes the object’s orientation and $\boldsymbol{p}^{o}_{t}\in\mathbb{R}^{3}$ the position. All simulation states have corresponding ground-truth values, denoted by the hat symbol. For instance, the reference object rotation is $\{\hat{\boldsymbol{\theta}}^{o}_{t}\}_{t=1}^{T}$. The environmental setup for the simulation is detailed in Sec. B.
@@ 3.1 Policy Representation
State. The state $\boldsymbol{s}_{t}$, which serves as input to the policy, comprises two components $\boldsymbol{s}_{t}=\{\boldsymbol{s}_{t}^{s},\boldsymbol{s}_{t}^{g}\}$. The first part, $\boldsymbol{s}_{t}^{s}$, contains human proprioception and object observations, expressed as, $\{\{\boldsymbol{\theta}_{t}^{h},\boldsymbol{p}_{t}^{h},\boldsymbol{\omega}_{t}^{h},\boldsymbol{v}_{t}^{h}\},\{\boldsymbol{\theta}_{t}^{o},\boldsymbol{p}_{t}^{o},\boldsymbol{\omega}_{t}^{o},\boldsymbol{v}_{t}^{o}\},\{\boldsymbol{d}_{t},\boldsymbol{c}_{t}\}\},$ where $\{\boldsymbol{\theta}_{t}^{h},\boldsymbol{p}_{t}^{h},\boldsymbol{\omega}_{t}^{h},\boldsymbol{v}_{t}^{h}\}$ represent the rotation, position, angular velocity, and velocity of all joints, respectively, while $\{{\boldsymbol{\theta}}_{t}^{o},\boldsymbol{p}_{t}^{o},{\boldsymbol{\omega}}_{t}^{o},\boldsymbol{v}_{t}^{o}\}$ represent the orientation, location, velocity, and angular velocity of the object, respectively. Motivated by [6], we include object geometry and whole-body haptic sensing from two elements: (i) $\boldsymbol{d}_{t}$, vectors from human joints to their nearest points on each object surface; and (ii) $\boldsymbol{c}_{t}$, contact markers indicating whether the human’s rigid body parts experience applied forces; this serves as simplified tactile or force sensing – an important multi-modal input in robot manipulation tasks [9]. The goal state $\boldsymbol{s}_{t}^{g}=\{\boldsymbol{s}_{t,t+k}^{g}\}_{k\in K}$ integrates reference poses from the ground truth motion, where $\boldsymbol{s}_{t,t+k}^{g}$ is defined as,
@@
Given that most MoCap data does not provide reference contact or tactile information, we extract reference contact markers $\hat{\boldsymbol{c}}_{t+k}$ by inferring dynamic information, beyond relying solely on inaccurate contact distances, specifically by analyzing the object’s acceleration to detect human-induced forces. To accommodate the variability in contact distances observed in reference motion, we discretize reference contact markers using varying distance thresholds, as illustrated in Fig. 3(i). The neutral areas serve as buffer zones, avoiding the penalization or enforcement of strict contact. See Sec. C of supplementary for details.
@@ 3.2 Imitation as Perfecting
Imitation as Retargeting. We tailor teacher policies to each human subject, while all policies share the same base human model. This serves the retargeting purpose by converting HOIs from different human shapes into a unified base shape. Although motion imitation does not necessarily require a unified human model [92, 51], our approach offers two benefits: (i) It enhances integration with kinematic generation methods, which generally perform better on a single, unified shape [17]. (ii) It demonstrates possible integration with real-world humanoid deployment, which requires retargeting to a consistent physical embodiment. In Figure LABEL:fig:teaser, our method translates MoCap data into motor skills on a Unitree G1 [81] with two Inspire hands [24], all without external retargeting in complex contact-rich scenarios. See Sec. F of the supplementary for additional details.
@@
Embodiment-Aware Reward. When the human and object are far apart, retargeting should prioritize capturing rotational motion, whereas when they are close, accurate position tracking becomes crucial for achieving contact. To reflect this, we define the weights $\boldsymbol{w}_{d}$ that are inversely proportional to the distances between joints and the object [113]. The reward thus includes cost functions for joint position $E_{p}^{h}=\langle\boldsymbol{\Delta}^{h}_{p},\boldsymbol{w}_{d}\rangle$, rotation $E_{\theta}^{h}=\langle\boldsymbol{\Delta}^{h}_{\theta},\boldsymbol{1}-\boldsymbol{w}_{d}\rangle$, and interaction tracking $E_{d}=\langle\boldsymbol{\Delta}_{d},\boldsymbol{w}_{d}\rangle$, where $\langle\cdot,\cdot\rangle$ is the inner product, $\boldsymbol{\Delta}^{h}_{p}[i]=\|\hat{\boldsymbol{p}}^{h}[i]-\boldsymbol{p}^{h}[i]\|$, $\boldsymbol{\Delta}^{h}_{\theta}[i]=\|\hat{\boldsymbol{\theta}}^{h}[i]\ominus\boldsymbol{\theta}^{h}[i]\|$, and $\boldsymbol{\Delta}_{d}[i]=\|\hat{\boldsymbol{d}}[i]-\boldsymbol{d}[i]\|$ represent the displacement for the variables defined in Sec. 3.1 with timestep $t$ omitted. The formulation of $\boldsymbol{w}_{d}$ is provided in Sec. D of supplementary. The reward to be maximized can be formulated as $\exp(-\lambda E)$ for each cost function $E$ with a specific hyperparameter $\lambda$. Details can be found in Sec. D.
@@
Physical State Initialization. Learning later-phase motion can be essential for policies to achieve high rewards during earlier phases, compared to incrementally learning from the starting phase. Thus, Reference State Initialization (RSI) [63] sets the current pose ${\boldsymbol{q}}_{t}$ to a reference pose $\hat{\boldsymbol{q}}_{t}$ at a random timestep $t$, for initializing the rollout. However, initializing with the imperfect reference can introduce critical artifacts, such as contact floating or incorrect hand motion, leading to unrecoverable failures, e.g., object falling, as depicted in Figure 3(ii). These issues render many initializations ineffective, limiting training on certain interaction phases since successful rollouts may not reach them before the maximum length. The problem is exacerbated by the use of prioritized sampling [92, 78], which favors high-failure-rate initializations.
@@ 3.3 Imitation with Distillation
Reference Distillation. Noisy MoCap data can hinder policy learning, especially at larger scales. In contrast, teacher policies trained on smaller-scale data effectively address these issues by correcting contact artifacts, refining hand placements, and recovering missing details (see Figures LABEL:fig:teaser and 5). To fully leverage teacher policies, we use their rollouts as references for defining the student policy’s goal state and reward functions, distinguishing our approach from distillation based on only action output.
Policy Distillation. We apply distillation on action outputs, which we view as crucial for scaling policies to large datasets. In essence, we trade space for time: teacher policies are trained in parallel on smaller data subsets, allowing the student policy to scale through distillation. Following Algorithm 1, we begin with Behavior Cloning (BC) [29] and use RL fine-tuning to go beyond demonstration memorization, an approach common in LLM alignment [56]. We integrate BC into online policy updates with a staged schedule: we start with DAgger [69] and gradually transition to PPO. Throughout, the critic is continuously trained with the reward from Sec. 3.2. This RL fine-tuning phase is crucial as teacher policies may behave differently when performing similar skills, and simple BC can lead to suboptimal “averaging” behavior, where RL fine-tuning helps the student converge on optimal solutions.